Amount ($million)

Weight

Cost of capital

Debt

US$ 26,162

0.00%

Equity

US$ 1,500

84.40%

Preferred stock

US$ 4,059

14.63%

Total

WACC
Answer #1
Calculation of WACC:
Value Weights(a) Return(b) (a*b)
Debt 26162 0.82 0 0.00
Equity 1500 0.05 84.4 3.99
Preferred stock 4059 0.13 14.63 1.87
31721 1 5.86
WACC is 5.86%
Calculation of weight:
Debt= 26162/31721=0.82
Equity=1500/31721= 0.05
Preferred stock=4059/31721=0.13